# Entanglement engineering of optomechanical systems by reinforcement learning

### v1

@@ III Deep reinforcement-learning control of quantum optomechanical systems with linear interactions
Figure 1: Workflow of measurement-based feedback control of deep RL to create and stabilize entanglement in an open quantum optomechanical system dissipatively coupled to the vacuum bath. (a) The RL agent acts in one or multiple parallel quantum optomechanical environments according to the parameterized policy and collects data in one episode consisting of $T$ time steps: observation $O_{t}$ of the photon number, reward $R_{t}=-|O_{t}-0.5|$, and action $G_{t}\in[-5,5]\omega_{m}$, after which the quantum optomechanical environment is reset. After one or several episodes, the policy is updated using minibatch data to maximize the accumulated reward. The aim is to achieve the desired entanglement between the cavity-optical and mechanical modes as best as possible. (b) An example: an entangled Bell state of the beam-splitter Hamiltonian or “swap” Hamiltonian. In the microwave-equivalent system of a quantum optomechanical system [82, 83, 84], the photon number statistics associated with the microwave cavity mode can be experimentally detected through weak measurements [69].
We consider a quantum optomechanical system in the regime where the interactions between the cavity and oscillator modes are linear, as shown in Fig. 1(a). Experimentally, directly measuring the entanglement, e.g., in terms of logarithmic negativity, for arbitrary entangled states is generally not feasible. It remains challenging to identify an experimentally feasible quantity to characterize the entanglement in arbitrary quantum systems. We focus on the relationship between logarithmic negativity and the expected photon number, based on recent experiments on homodyne detection of the photon number [69, 59, 45, 55, 57].
III.1 PPO RL agent as applied to quantum optomechanical systems
A quantum optomechanical system with linear photon-phonon interactions is governed by the beam-splitter Hamiltonian in Eq. (3). In an optical experimental platform, a 50:50 beam splitter with the transformation angle $\pi/4$ can create an entangled Bell state between the two input optical modes [85, 86, 87]. Similarly, in a quantum optomechanical system, Bell states between photon and phonon modes can be realized by controlling the beam-splitter Hamiltonian. The maximally attainable value of the logarithmic negativity is $E_{N}\sim\log 2$, corresponding to the maximally entangled Bell state shown in Fig. 1(b). This “best” entangled state can be realized by the PPO agent, regardless of whether the observation is the expectation or random measurement of the photon number. To see this, we note that, in the beam-splitter model, the initial quantum state is set as a pure state [88, 89]
$|\psi\rangle=|10\rangle,$ (8)
where the photon is in the first excited mode and the phonon is in the vacuum mode. A mixed state is assumed to be in the form of [90, 57]
$\rho=(1-p)|10\rangle\langle 10|+p|01\rangle\langle 01|,$ (9)
with the probability $p$ being a random number between zero and one because of the coupling to the classical environment. The partial observable of the quantum state is set as the expectation of the photon number $\langle\hat{n}_{p}\rangle(t)$ or the WCM flow
$\mathcal{I}(t)=\langle\sqrt{\eta}\hat{P}_{1}\rangle(t)+\frac{dW(t)}{\sqrt{4\eta}\;dt},$ (10)
where $\eta$ is the measurement rate, $\hat{P}_{1}=|1\rangle\langle 1|$ is the measurement projector on the Fock state $|1\rangle$, and $dW(t)$ is the Wiener increment with zero mean and variance $dt$. To proceed, we note that the beam-splitter Hamiltonian is limited to a four-level basis, due to the following reasons: (1) only one energy level has been excited from the vacuum state - a cavity mode or an oscillator mode, (2) the system couples to the vacuum bath only at absolute zero temperature (i.e., without any thermal excitation), and (3) the linear interaction serves only to transfer the quantum states between the cavity and mechanical mode (i.e., no quantum excitation), where any interactions between higher-level quantum states are blocked. In this case, the maximum logarithmic negativity $E_{N}\sim\log 2$ (in natural base) implies that the attained quantum state is the following Bell state, as illustrated in Fig. 1(b):
$|\Phi^{\varphi}\rangle=\frac{1}{\sqrt{2}}\left[|10\rangle+e^{i\varphi}|01\rangle\right],$ (11)
with the associated expected photon number $\langle\hat{n}_{p}\rangle=\langle\hat{P}_{1}\rangle=0.5$.
The first term in the weak measurement flow in Eq. (10), which contains no noise, is related to the expectation of the photon number. As a result, the reward function can be set as
$R_{t}\equiv-|O_{t}-0.5|,$ (12)
regardless of whether the observation $O_{t}$ is $\langle\hat{n}_{p}\rangle(t)$ or $\mathcal{I}(t)$. Note that the beam-splitter Hamiltonian transforms the initial state $|10\rangle$ or $|01\rangle$ to a Bell state with the corresponding expected photon number: $\langle\hat{P}_{1}\rangle=0.5$, where the strength of the coupling to the environment is much weaker than the beam-splitter interaction. However, if the initial state is the mixed state in Eq. (9), the $|10\rangle\langle 10|$ and $|01\rangle\langle 01|$ components will become independently entangled, resulting in the total quantum state being a mixture of two entangled Bell states. As a result, a nontrivial entanglement value is expected for the initial mixed state governed by the beam-splitter Hamiltonian. Because of the relatively small target value of the expected photon number: $\langle\hat{P}_{1}\rangle=0.5$, the variance $\mathcal{I}(t)$ in the WCM flow can be reduced by a Gaussian filter [91] with the weak measurement rate $\eta\leq 1$. The Gaussian kernel parameters of the filter such as the filter interval and the variance can be numerically chosen to reduce the standard deviation of the measurement current into a certain range, e.g., about ten times larger than the mean value (see Appendix E for details). The PPO agent applies an updated stochastic policy to the quantum optomechanical environment to maximize the accumulated reward, where the action $G(t)$ is proportional to the amplitude of the cavity mode: $G(t)=g_{0}\bar{\alpha}_{c}(t)$. The action can be controlled by an incident laser [68] and is continuous in a certain range, e.g., $G\in[-5,5]\omega_{m}$.
As illustrated in Fig. 1, during online training, given a fixed training episode length, e.g., ${\rm Episode}=3000$, the RL agent bootstraps itself by executing the procedure described in Appendix D.1. In the initial preparation process, $\mathbb{N}$ identical and independent quantum optomechanical environments ($\mathbb{N}$ parallel environments) are prepared, where $\mathbb{N}=5$. In addition, the agent, which has two independent neural networks: actor and critic, is also initialized. The initial quantum state is based on Eq. (8) or Eq. (9) and the quantum environments are governed by the SME in Eq. (II.2).
In episodic learning, the quantum environments are reset after each episode. For each set of $\mathbb{Z}$ episodes (e.g., $\mathbb{Z}=5$), the agent obtains the observation $O_{t}$ about the photon number and the reward value $R_{t}=-|O_{t}-0.5|$ in Eq. (12) from $\mathbb{N}$ quantum environments, and independently acts on them by the current stochastic policy $\pi(G_{t}|O_{t};\boldsymbol{\theta})$. Essentially, the policy is the conditional probability distribution on the action space $G\in[-5,5]\omega_{m}$ given the observation $O_{t}$ and is parameterized through $\mathcal{\boldsymbol{\theta}}$. The $\mathbb{N}\times\mathbb{Z}$ independent trajectories, denoted as $\tau^{j}$ with the trajectory index $j=1,2,\ldots,\mathbb{N}\times\mathbb{Z}$, are collected with length $T=500$ (the number of time steps for each episode) and the step size $dt=0.01\omega_{m}^{-1}$. Each trajectory $\tau^{j}$ is a sequence of states (observations), actions, rewards, and next states (next observations):
$\tau^{j}=(O^{j}_{0},G^{j}_{0},R^{j}_{0},O^{j}_{1},\ldots,O^{j}_{T-1},G^{j}_{T-1},R^{j}_{T-1}),$ (13)
which can be organized as a sub-trajectory tuple
$\tau^{j}_{t}=(O^{j}_{t},G^{j}_{t},R^{j}_{t},O^{j}_{t+1})$ (14)
with the time stage index $t=0,1,\ldots,T-2$. At the terminal stage $t=T-1$, we have
$\tau^{j}_{T-1}=(O^{j}_{T-1},G^{j}_{T-1},R^{j}_{T-1}).$ (15)
For each sub-trajectory tuple $\tau^{j}_{t}$, the generalized advantage estimation (GAE) [92] $\hat{A}^{j}_{t}$ uses a value function estimator:
$\hat{A}^{j}_{t}=\delta^{j}_{t}+(\gamma\lambda)\delta^{j}_{t+1}+\ldots+(\gamma\lambda)^{T-t-1}\delta^{j}_{T-1}$ (16)
with
$\delta^{j}_{t}=R^{j}_{t}+\gamma V(O^{j}_{t+1};\boldsymbol{\phi})-V(O^{j}_{t};\boldsymbol{\phi}),$ (17)
where the value function $V(O^{j}_{t};\boldsymbol{\phi})$ is utilized to score the quality of $O^{j}_{t}$ based on the accumulated reward and parameterized by $\boldsymbol{\phi}$ and $\delta^{j}_{t}$ is the relative advantage of the current action selected by the stochastic policy $\pi(G^{j}_{t}|O^{j}_{t};\boldsymbol{\theta})$ with the discounted factor $\gamma\in(0,1)$ and hyperparameter $\lambda$ with typical value $\lambda=0.95$. Intuitively, $\hat{A}^{j}_{t}$ is utilized to numerically quantify the relative cumulative advantage of a certain action selected by the current stochastic policy from time $t$ to the terminal stage $T-1$, in which the future impact is included but regarded as less important than the corresponding previous one by the discount factor $\gamma\in(0,1)$. The finite-horizon discounted return $\hat{\mathcal{G}}^{j}_{t}$ is defined as
$\hat{\mathcal{G}}^{j}_{t}=\sum_{k=t}^{T-1}\gamma^{k-t}R^{j}_{k},$ (18)
which can be also obtained from the generalized advantage by
$\hat{\mathcal{G}}^{j}_{t}=\hat{A}^{j}_{t}+V(O^{j}_{t};\boldsymbol{\phi}),$ (19)
where $\hat{\mathcal{G}}^{j}_{t}$ denotes the accumulated reward from time $t$ to the terminal stage in the discounted version.
The neural networks constituting the actor and critic are updated from minibatches with size $\mathbb{M}$ from ××$\mathbb{N}\times\mathbb{Z}\times T$ data points, consisting of the sub-trajectory $\tau^{j}_{t}$, the generalized advantage $\hat{A}^{j}_{t}$ and the return $\hat{\mathcal{G}}^{j}_{t}$ over $k=10$ epochs with the Adam algorithm. The typical batch size is $\mathbb{M}=\mbox{int}(\mathbb{N}\times\mathbb{Z}\times T/10)$. For each epoch, the critic parameters $\boldsymbol{\phi}$ in the loss $L_{critic}(\boldsymbol{\phi})$ and the actor parameters $\boldsymbol{\theta}$ in the loss $L_{actor}(\boldsymbol{\theta})$ need to be updated to minimize the loss function over a random minibatch data. The mean square loss $L_{critic}(\boldsymbol{\phi})$ about the target $\hat{\mathcal{G}}_{i}$ for the value function $V(O_{i};\boldsymbol{\phi})$ is
$L_{critic}(\boldsymbol{\phi})=\hat{\mathbb{E}}_{i}[(V(O_{i};\boldsymbol{\phi})-\hat{\mathcal{G}}_{i})^{2}]$ (20)
and the clipped loss $L_{actor}(\boldsymbol{\theta})$ is given by
$L_{actor}(\boldsymbol{\theta})=\hat{\mathbb{E}}_{i}\left[-\textnormal{min}(r_{i}(\boldsymbol{\theta})\hat{A}_{i},\textnormal{clip}(r_{i}(\boldsymbol{\theta}),[1-\epsilon,1+\epsilon])\hat{A}_{i})\right],$ (21)
where $\hat{\mathbb{E}}_{i}[\;\;]=\sum_{i=1}^{\mathbb{M}}[\;\;]_{i}/\mathbb{M}$ is the empirical average over a minibatch of the data and $[\;\;]_{i}$ denotes the $i$th element of the minibatch with $i=0,1,\ldots,\mathbb{M}-1$, and the clip function $\textnormal{clip}(x,[\textnormal{min},\textnormal{max}])$ returns $x$ clipped to set limits: $\textnormal{min}\leq x\leq\textnormal{max}$. The probability ratio $r_{i}(\boldsymbol{\theta})>0$ between the current and old policies is
$r_{i}(\boldsymbol{\theta})=\frac{\pi_{\boldsymbol{\theta}}(G_{i}|O_{i})}{\pi_{\boldsymbol{\theta}_{\textnormal{old}}}(G_{i}|O_{i})}.$ (22)
If the current policy is the same as the old policy, we have $r_{i}(\boldsymbol{\theta}_{\textnormal{old}})=1$. In general, the ratio $r_{i}(\boldsymbol{\theta})$ needs to be away from the value one for the policy to be optimized. However, $r_{i}(\boldsymbol{\theta})$’s deviating too much from the value one will result in many fast policy updates, possibly leading to instabilities and even a collapse of the learning process. To avoid this, the clip function in the actor loss $L_{actor}(\boldsymbol{\theta})$ can be utilized to remove the incentive for $r_{i}(\boldsymbol{\theta})$ outside of the interval $[1-\epsilon,1+\epsilon]$ with typical clip range $\epsilon=0.2$, which decreases the updating speed of policy and improves the learning stability.
Intuitively, the goal of RL is to maximize the cumulative reward. In the case illustrated in Fig. 1, the objective is to achieve the entangled Bell state as fast as possible or, as stipulated by the reward function, to achieve the optimal photon number $O_{t}\rightarrow 0.5$ and to maintain this for as long as possible. When the RL agent converges to the optimal policy, the Bellman equation is satisfied [35], so the optimal value function satisfies
$V^{*}(O^{j}_{t};\boldsymbol{\phi})=R^{j}_{t}+\gamma V^{*}(O^{j}_{t+1};\boldsymbol{\phi}),$ (23)
i.e., the optimal value function about $O^{j}_{t}$ is equal to the current reward plus the future discounted cumulative reward, in which $O^{j}_{t+1}$ is determined by the action selected by the optimal policy $\pi^{*}(G^{j}_{t}|O^{j}_{t};\boldsymbol{\theta})$. It guarantees that the agent makes the best possible decisions to maximize the rewards [35]. Moreover, under the optimal policy, it means the zero generalized advantage $\hat{A}^{j}_{t}$, so the zero actor loss $L^{*}_{actor}$ is obtained. It is worth noting that the optimal value function is equal to the discounted accumulated reward from Eq. (18):
$V^{*}(O_{i};\boldsymbol{\phi})=\hat{\mathcal{G}}_{i},$ (24)
which also gives zero critic loss $L_{critic}(\boldsymbol{\phi})$. In the online training process, the RL agent trained as described is called the PPO agent, whose policy is randomly initialized and will gradually converge to the optimal one under the described training scenarios to achieve the maximum accumulated reward. Physically, this enables the entangled Bell state to be created and stabilized. For online testing, the optimized policy is no longer updated and only one quantum environment is involved.
III.2 Superiority of deep RL-based control of entanglement engineering for benchmark control methods

### v2

@@ Introduction
Figure 1: Experimental proposal of measurement-based feedback control of deep RL to create and stabilize entanglement in an open quantum optomechanical system dissipatively coupled to the vacuum bath. Quantum optomechanics was experimentally realized in a microwave electromechanical system [58, 59, 60], where the multiplexing qubit was used to weakly couple to the microwave resonator for extracting the photon number statistics through weak measurements [49]. The RL agent acts in one or multiple parallel quantum optomechanical environments according to the parameterized policy and collects data in one episode consisting of $T$ time steps: observations $O_{t}$, reward $R_{t}$, and actions, after which the quantum optomechanical environment is reset. After one or several episodes, the policy is updated using minibatch data to maximize the accumulated reward. The aim is to achieve the desired entanglement $E_{N}\sim\log 2\sim 0.7$ (in the natural logarithmic base) between the cavity-optical and mechanical modes. Entanglement engineering of this type can be achieved in both the linear and nonlinear interaction regimes. In the linear case, the task is similar to that of achieving an entangled Bell state of the beam-splitter Hamiltonian or “swap” Hamiltonian. In the nonlinear regime, the entangled states from entanglement engineering can be complicated. Illustrated are the resulting photon and phonon number distributions of the entangled states.
Results
Experimental proposal for entanglement engineering
Our goal is achieving entanglement engineering between the optical cavity and mechanical oscillator modes using deep RL. Based on the current experimental progress, we articulate an experimental proposal to achieve this goal, as shown in Fig. 1. Consider a Fabry-Perot cavity that consists of a single-mode cavity and a movable end mirror. The optical cavity has the frequency $\omega_{c}$ and the optical field exerts a radiation pressure on the mirror. The cavity mode is driven externally by a coherent laser field with frequency $\omega_{L}$. The mirror’s quantized center-of-mass motion is described by a harmonic oscillator of frequency $\omega_{m}$. In the rotating frame of the laser, the Hamiltonian describing the coupling between the optical cavity and mechanical oscillator modes is given by [55, 56]
$\tilde{H}_{nl}=-\hbar\Delta\hat{a}^{\dagger}\hat{a}+\hbar\omega_{m}\hat{b}^{\dagger}\hat{b}+\hbar g_{0}(\hat{b}^{\dagger}+\hat{b})\hat{a}^{\dagger}\hat{a}+\hbar\alpha_{L}(\hat{a}^{\dagger}+\hat{a}),$
where $\hat{a}$ and $\hat{b}$ are the annihilation operators of the cavity and mechanical mode, respectively, $\hat{a}^{\dagger}$ and $\hat{b}^{\dagger}$ are the corresponding creation operators. The frequency detuning of the cavity is $\Delta\equiv\omega_{L}-\omega_{c}$. The nonlinear coupling $g_{0}$ arises from the radiation pressure force between the light and the movable mirror (details given in Supplementary Note 2), and $\alpha_{L}$ is the real amplitude of the driven electromagnetic field. We set $g_{0}>\kappa$ so that the single-photon optomechanical coupling rate $g_{0}$ exceeds the coupling strength $\kappa$ between the cavity and the vacuum bath. This condition guarantees observable nonlinear quantum effects [61]. Under the strong laser approximation: $|\bar{\alpha}_{c}|\gg 1$, where $|\bar{\alpha}_{c}|$ is the amplitude of the light field inside the cavity induced by the strong laser, we have $\hat{a}\approx\bar{\alpha}_{c}+\delta\hat{a}$ with $\delta\hat{a}$ denoting the excitation or the shifted oscillator on top of the large coherent state with the amplitude $\bar{\alpha}_{c}$. The resulting linearized beam-splitter or “swap” Hamiltonian  [62, 55] is
$\tilde{H}_{bs}\approx\hbar\omega_{m}\delta\hat{a}^{\dagger}\delta\hat{a}+\hbar\omega_{m}\hat{b}^{\dagger}\hat{b}+\hbar G(\delta\hat{a}^{\dagger}\hat{b}+\hat{b}^{\dagger}\delta\hat{a}),$
which is obtained in the red-detuned regime $\Delta=-\omega_{m}$, where the coefficient $G\equiv g_{0}\bar{\alpha}_{c}$ can be tuned by the amplitude of the incoming laser (a time-dependent modulation) [63]. The interaction term describes the state transfer between photons and phonons in the strong coupling regime for $G>\kappa$, with $\kappa$ ($\gamma$) being the decay rate of the cavity (mechanical) mode to the vacuum bath at zero temperature.
Our control strategy was developed based on considering the current experimental capability. Previous works on the microwave regime of the optomechanical systems [58, 59, 60] suggested the feasibility of the experimental implementation of our RL control scheme. In particular, a one-to-one correspondence between the Fabry-Perot cavity and the microwave electromechanical system was demonstrated [64, 59, 58]. As shown in Fig. 1, the microwave resonator of an LC circuit is equivalent to the Fabry-Perot optical cavity mode with the movable capacity [64] $C_{m}(x)$ corresponding to the flexible mirror in the optical cavity. The resistors $R_{c}$ and $R_{m}$ can be related to the decay rate $\kappa,\gamma$ to the vacuum bath [64]. Based on the experimental results, we can compare the typical parameter configurations between the optomechanical and electromechanical systems. The decay rate of the optical cavity mode is $\kappa=0.01\,\omega_{m}$ in the linear regime and $\kappa=0.1\,\omega_{m}$ in the nonlinear regime, with the better quality of the mechanical oscillator mode $\gamma=0.01\,\kappa$. Consequently, we have $\gamma\approx 10^{-3}\,\omega_{m}\sim 10^{-4}\,\omega_{m}$. The typical experimental decay rate of the microwave resonator is [59, 58, 65, 66, 67, 61] $\kappa\approx 0.01\,\omega_{m}\sim 0.1\,\omega_{m}$ with $\gamma\approx 10^{-3}\,\omega_{m}\sim 10^{-9}\,\omega_{m}$. In our work, the nonlinear coupling is set to be $g_{0}=0.2\,\omega_{m}$, whereas the typical coupling in the strong coupling regime in a previous work [58] was about $g_{0}=0.1\,\omega_{m}$. The strength of the laser in our work is $G\in[-5,5]\,\omega_{m}$ for the linear system in the red-detuned regime $\Delta=-\omega_{m}$ and $\Delta,\alpha_{L}\in[-5,5]\,\omega_{m}$ for the nonlinear system. In the microwave version, this range can be adjusted by the pump’s strength [59, 58, 65, 66, 67, 61].
In the microwave regime, it was demonstrated that the photon-number statistics of a microwave cavity mode can be detected using multiplexed photon number measurements [49, 68, 69]. By this method, the multiplexing qubit encodes multiple bits about the photon number distribution of a microwave resonator through dispersive interaction. A frequency comb drive, distributed at $f_{\textnormal{MP}}-k\chi$, reads out all the information about the photon number distribution at once [49], where $k$ denotes the number of photons and $\chi$ represents the dispersive qubit-resonator coupling, as shown in Fig. 1. The reduction in the reflection amplitude, $1-r_{k}$ with $k=0,1,...$, of the frequency comb, is proportional to the photon-number distribution of the microwave cavity mode over the Fock bases, as detected by the weak measurement [29, 49]. In our circuit design of experimental proposal, we add a capacitor $C_{q}$ to realize the weak coupling to the original electromechanical system. The coupling capacitance is small enough to be neglected in the total Hamiltonian, but it still allows the multiplexing qubit, denoted by the green cross in Fig. 1, to encode the photon number distribution of the microwave resonator through dispersive interaction.
Under weak measurement [29, 49], the sequence of the reduced reflection amplitude $1-r_{k}$ is collected by the PPO agent, which is proportional to the occupied photon number probability. Consequently, the expected photon number is calculated as
$\langle\hat{n}_{p}\rangle=\sum_{n}n\langle\sqrt{\eta}\hat{P}_{n}\rangle/\sqrt{\eta}=\sum_{n}n\langle\hat{P}_{n}\rangle$
and the WCM photocurrent is
$\sqrt{\eta}\,\mathcal{I}(t)=\sum_{n}n\left[\langle\sqrt{\eta}\hat{P}_{n}\rangle+\frac{dW_{n}(t)}{\sqrt{4\eta}dt}\right]$ (1)
with the measurement rate $\eta$, where $\hat{P}_{n}=|n\rangle\langle n|$ is the measurement projector on the Fock state $|n\rangle$, and $dW(t)$ is the Wiener increment with zero mean and variance $dt=0.01\,\omega_{m}^{-1}$ (the time step size in our calculations). In the linear quantum optomechanical regime, the Fock space for each mode is limited to $n=0,1$. The action is the amplitude modulation of the laser, which is in the range $G\in[-5,5]\,\omega_{m}$. In the nonlinear regime, the Fock dimension is $n=0,1,\ldots,9$. The time-dependent control signal consists of the detuning $\Delta$ and the amplitude $\alpha_{L}$ of the driven laser within the fixed range $\Delta,\alpha_{L}\in[-5,5]\,\omega_{m}$.
The open dissipative quantum optomechanics under the WCM obey the stochastic master equation (SME) (see Methods). The number $n_{\textnormal{traj}}$ of trajectories simulated from SME can be selected according to the following considerations. If the observable is some expected physical quantity, using one trajectory is sufficient to extract the information about the quantum state: $n_{\textnormal{traj}}=1$. Experimentally, WCMs are performed, encoding the Wiener process in the observation and resulting in a large variance from the expectation value. To reduce the variance, more quantum trajectories should be used. To make computations feasible, we use five trajectories: $n_{\textnormal{traj}}=5$.
In the online training phase, for each episode with time steps, e.g., $T=500$, the PPO agent - the combination of the actor and critic network, collects the sequence of the observations $O(t)=\langle\hat{n}_{p}\rangle(t)$ or $\mathcal{I}(t)$, the reward value $R(t)=-|O(t)-\langle\hat{n}^{\textnormal{target}}_{p}\rangle(t)|$, and the resulting actions generated by its policy. After one or several episodes, the policy of the PPO agent is updated using minibatch data to maximize the accumulated reward. The RL agent is designed to interact with a single or multiple parallel quantum environments to make the time evolving observation $O(t)$ align with the target one $\langle\hat{n}^{\textnormal{target}}_{p}\rangle(t)$. In the online testing phase, the policy of the well-trained agent will not update and only interact with a single quantum environment to give the optimal control protocol to the corresponding observation. To realize entanglement engineering, i.e., achieving the desired entanglement between the cavity-optical and mechanical modes, finding the relation between the experimental observables and entanglement quantities is an unavoidable challenge. In our work, the model-free PPO agent finds the numerical relationship between them and realizes the entanglement engineering in both the linear and nonlinear regimes of quantum optomechanics, as shown in Fig. 1.
A general quantity to measure the entanglement between arbitrary quantum bipartite systems for any mixed states is the logarithmic negativity [70, 71, 72], without the influence of the vacuum bath [73]. In contrast, the conventional pure-state entanglement measures, such as the von Neumann and Rényi entropy, capture both quantum and classical correlations. Since the goal of our study is harnessing the entanglement between the cavity and oscillator modes, we focus on the logarithmic negativity: $E_{N}(\rho)\equiv\log_{2}||\rho^{T_{i}}||_{1}$, where $||X||_{1}=\textnormal{Tr}\sqrt{X^{\dagger}X}$ is the trace norm of the partial transpose $\rho^{T_{i}}$ with respect to the two subsystems $i=0$ (quantum-optical cavity mode) and $1$ (mechanical oscillator mode). The logarithmic negativity measures the degree to which $\rho^{T_{i}}$ fails to be positive, i.e., the extent of inseparability or entanglement, and it is the upper bound of the distillable entanglement [70, 71]. The logarithmic negativity is the full entanglement monotone [71], which satisfies the following criteria [74, 72]: (1) $E_{N}$ is a non-negative functional, (2) $E_{N}$ vanishes if the state $\rho$ is separable, and (3) $E_{N}$ does not increase on average under Gaussian local operations and classical communication [75, 76] or positive partial transpose preserving operations [77]. Since $E_{N}$ quantifies the quantum correlation between the bipartite systems in spite of the coupling to the vacuum bath, the value of $E_{N}$ calculated from the cavity mode is equal to that of the oscillator mode: $E^{0}_{N}=E^{1}_{N}=E_{N}$, which can be verified numerically.